# Landau-Zener-Stückelberg interference in edge state pumping

### v1

@@ Abstract
Conclusion: In this work we interpret the ESP in terms of LZS interference, breaking the adiabatic condition at the two NAPs. This LZS mechanism naturally explains the oscillation of ESP as a function of evolution time $T$. In the disordered case, we show that the possible anticrossing between the edge state and the bulk state may also lead to failure of ESP. Our results show that the ESP realized in the current experiments need to be examined in much longer chains under more stringent conditions Rechtsman et al. (2013); Zhang et al. (2018); Jürgensen et al. (2021); Chen et al. (2021). This work puts forward an open question how one preserves the adiabatic condition during ESP by engineering of the driven protocol to achieve perfect adiabatic transport in a much shorter time Privitera et al. (2018); Citro and Aidelsburger (2023); Longhi (2019), which is essential for the manipulation of topological edge states in large, realistic systems. These results can be simulated using ultracold atoms Schomerus (2013); Viebahn et al. (2019); Nakajima et al. (2021), optical and acoustic waveguides Xia et al. (2021); Wang et al. (2023a) and electric circuits Li et al. (2023); Mei et al. (2018), in which the nonlinearity and non-Hermitian Jürgensen et al. (2021); Lin et al. (2022); Wang et al. (2018) can be naturally included. The new platform for realization of LZS in topological models with different types of symmetries Qi and Zhang (2011); Hasan and Kane (2010); Schnyder et al. (2008) will also broaden the family of this intriguing dynamics and unfold the associated nonadiabatic physics.
@@
Acknowledgments: Y. L. is supported by Young Scientist Fund [NSFC11804087], Natural National Science Foundation [NSFC12047501, NSFC12074107]; and Science and Technology Department of Hubei Province [2022CFB553, 2022CFA012, 2024AFA038].

### v2

@@ Abstract
Conclusion: In this work we interpret the ESP in terms of LZS interference, which countertuitively breaks the adiabatic condition at two NAPs. This LZS interference at NAPs naturally explains the oscillation of ESP as a function of evolution time, and also contributes to its efficient pumping. In the disordered case, we reveal that the possible anticrossing between the edge and the bulk states even leads to failure of ESP. Our results show that the ESP available in the current experiments requires scrutinisation in much longer chains under more stringent conditions Rechtsman et al. (2013); Zhang et al. (2018); Jürgensen et al. (2021); Chen et al. (2021). We put forward an open question how one preserves the adiabatic condition during ESP, by engineering the driven protocol to achieve transport in a much shorter time Privitera et al. (2018); Citro and Aidelsburger (2023); Longhi (2019), which is essential for the manipulation of topological edge states in large, realistic systems. Our theoretical predictions can be simulated using ultracold atoms Schomerus (2013); Viebahn et al. (2019); Nakajima et al. (2021), optical and acoustic waveguides Xia et al. (2021); Wang et al. (2023a) and electric circuits Li et al. (2023); Mei et al. (2018), in which the nonlinearity and non-Hermitian Jürgensen et al. (2021); Lin et al. (2022); Wang et al. (2018) can be naturally included. This unique platform for realizing of LZS model in our work, when using the various topological models with different symmetries Qi and Zhang (2011); Hasan and Kane (2010); Schnyder et al. (2008), will also greatly broaden the family of LZS systems for this intriguing dynamics, and unfold their associated nonadiabatic physics Ivakhnenko et al. (2023).
@@
Acknowledgments: Y. L. is supported by Young Scientist Fund [NSFC11804087], Natural National Science Foundation [NSFC12047501, 12074107 and 12075081]; and Science and Technology Department of Hubei Province [2022CFB553, 2022CFA012, 2024AFA038].